### v1

@@ 3.2 From Vertex Cover to Diverse Vertex Cover
Now we can define dynamic programming entries representing the tuples of partial solutions for each node $t$ of the decomposition. Intuitively, these entries should be combinations of entries from $\mathcal{F}_{t}$. For $t\in V(\mathcal{D})$ let $\mathcal{R}_{t}$ be a set of tuples
$\ell=max\{Div(S_{1},\ldots,S_{r})|\forall_{j\in[r]}\forall_{i\in[\omega]}S_{j}\textrm{ is a vertex cover of }G_{t}\textrm{ that uses }c_{i}^{j}\textrm{ vertices of color }i\}.$
Note that $|\mathcal{R}_{t}|\leq(k+1)^{r\omega}$.
  1. 1.
    $t=intro(v,a)$
    $\mathcal{R}_{t}=\{(C^{1},\ldots,C^{r},\ell)|\forall_{j\in[r]}C^{j}\in\{C_{\{v\}},C_{\emptyset}\},\ell=f(m(C^{1},\ldots,C^{r},v))\},$
    $\mathcal{R}_{t}$ can be computed in time $2^{r}\cdot r\omega$.
  2. 2.
    $t=t_{1}\oplus t_{t}$
    $\mathcal{R}_{t}=\{(C^{1},\ldots,C^{r},\ell)|\exists_{(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})\in\mathcal{R}_{t_{1}},(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})\in\mathcal{R}_{t_{2}}}\forall_{j\in[r]}C^{j}=C_{1}^{j}+C_{2}^{j}\land|C^{j}|\leq k,\\ \ell=max\{\ell_{1}+\ell_{2}-|V(G)|f(0^{r})|\exists_{(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})\in\mathcal{R}_{t_{1}},(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})\in\mathcal{R}_{t_{2}}}\ s.t.\ \forall_{j\in[r]}C^{j}=C_{1}^{j}+C_{2}^{j}\land|C^{j}|\leq k\}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{2r\omega}\cdot r\omega$.
    Since $V(G_{t_{1}})$ and $V(G_{t_{2}})$ are disjoint, $(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})$ represents a tuple $S_{1},\ldots S_{r}\subseteq V(G_{t_{1}})$ of diversity $\ell_{1}$ and $(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})$ represents a tuple $P_{1},\ldots P_{r}\subseteq V(G_{t_{2}})$ of diversity $l_{2}$, by Lemma 3.2 the diversity of $S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r}$ represented by $(C^{1},\ldots,C^{r},\ell)$ is $\ell_{1}+\ell_{2}-|V(G)|\cdot f(0^{r})$. Observe that $(C^{1},\ldots,C^{r})$ may represent multiple tuples of partial solutions in $G_{t}$, but since we are interested in maximizing the diversity, it suffices to keep only the best diversity value among all these tuples.
  3. 3.
    $t=recolor(t^{\prime},a\rightarrow b)$
    $\mathcal{R}_{t}=\{(\rho_{a\rightarrow b}(C^{1}),\ldots,\rho_{a\rightarrow b}(C^{r}),\ell),|(C^{1},\ldots,C^{r},\ell)\in\mathcal{R}_{t^{\prime}}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{r\omega}\cdot 2r$.
  4. 4.
    $t=addEdges(t^{\prime},a,b)$
    $\mathcal{R}_{t}=\{((c_{1}^{1},\ldots,c_{\omega}^{1}),\ldots,(c_{1}^{r},\ldots,c_{\omega}^{r}),\ell)\in\mathcal{R}_{t^{\prime}}|\forall{j\in[r]}c_{a}^{j}=n_{a}\lor c_{b}^{j}=n_{b}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{r\omega}\cdot 2r$.
Theorem 3.3.
$\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2r\omega}\cdot r\omega).$
Proof.
For each $t\in V(\mathcal{D})$ $\mathcal{R}_{t}$ can be computed in time
4 Monotone DP for Vertex Problems
In the Vertex Cover example each partial solution for a node $t$ of the decomposition was a disjoint union of partial solutions for the children of $t$. That is why we could easily leverage this dynamic programming to the diverse version, using Lemma 3.2 to compute the diversity. It tuns out that this property holds for many dynamic programming algorithms for vertex problems on cliquewidth decompositions. We will call such dynamic programming monotone. In this chapter we define formally what is a dynamic programming and monotone property and prove that having a monotone dynamic programming for each of the vertex problems $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ and a cliquewidth decomposition $\mathcal{D}$ of $G$, we can easily solve $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on $G$.
4.1 Definitions
Definition 4.1 (Dynamic programming core).
A dynamic programming core is an algorithm $\mathfrak{C}$ that takes cliquewidth decomposition $\mathcal{D}$ of a graph $G$ and produces:
  • A finite set $Accept_{\mathfrak{C},\mathcal{D}}\subseteq\{0,1\}^{*}$,
Let $\tau(\mathfrak{C},\mathcal{D})$ be the time necessary to construct these data and $Size_{\mathfrak{C},\mathcal{D}}=max_{t\in V(\mathcal{D})}|Process_{\mathfrak{C,\mathcal{D}}}(t)|$.
Intuitively, $Process_{\mathfrak{C},\mathcal{D}}$ corresponds to the transitions of the dynamic programming. $ww_{1}\ldots w_{\delta(t)}\in Process_{\mathfrak{C},\mathcal{D}}(t)$ if a partial solution represented by $w$ at node $t$ can be obtained from partial solutions represented by $w_{1}\ldots w_{\delta(t)}$ at its children $t_{1}\ldots t_{\delta(t)}$.
Definition 4.2 (Witness).
  • for each $t\in V(\mathcal{D})$ with children $t_{1},\ldots,t_{\delta(t)}$ $\alpha(t)\alpha(t_{1})\ldots\alpha(t_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)$,
  • $\alpha(q)=w$.
Definition 4.3.
We say that a dynamic programming core $\mathfrak{C}$ solves a graph problem $\mathcal{P}$ if for each graph $G$ for each cliquewidth decomposition $\mathcal{D}$ of $G$, $G\in\mathcal{P}$ if and only if a $(\mathfrak{C},\mathcal{D})$-witness exists.
Theorem 4.4.
Proof.
  • If $t$ has children $t_{1},\ldots t_{\delta(t)}$, $\Pi(t)=\{w|\exists_{(w,w_{1},\ldots w_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)}\forall_{i\in[\delta(t)]}w_{i}\in\Pi(t_{i})\}$.
$G\in\mathcal{P}$ if and only if $\Pi(q)\cap Accept_{\mathfrak{C},\mathcal{D}}\not=\emptyset$, where $q$ is the root of $\mathcal{D}$. ∎
Definition 4.5 (Monotonicity).
Let $\mathcal{P}$ be a vertex problem. We call a dynamic programming core $\mathfrak{C}$ monotone if there exists a function $\rho:\{0,1\}^{*}\rightarrow\{0,1\}$ such that for each graph $G$, for each cliquewidth decomposition $\mathcal{D}$ of $G$, for each $S\subseteq V(G)$, $(G,S)\in\mathcal{P}$ if and only if there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that ${S=\{v\in V(G):\rho(\alpha(Leaf_{\mathcal{D}}(v)))=1\}}$. We call such $\rho$ a $\mathfrak{C}$-vertex-membership function and say the pair $(\mathfrak{C},\rho)$ solves $\mathcal{P}$.
Intuitively, a vertex membership function $\rho$ together with a cliquewidth decomposition $\mathcal{D}$ of a graph $G$ and a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ provide an encoding of a subset of vertices of $G$, using only the leaves of $\mathcal{D}$.
Let $S_{\rho}(\mathcal{D},\alpha)=\{v\in V(G):\rho(\alpha(Leaf_{\mathcal{D}}(v)))=1\}$. Note that in this definition we determine whether a vertex is in the subset or not based only on the value of the witness on its introduce node.
4.2 Proof of Theorem 1.2
The following theorem is a formal reformulation of Theorem 1.2 formulated in the introduction. Indeed, if $t_{\mathcal{D}}=\mathcal{O}(\tau(\mathfrak{C}_{i},\mathcal{D})+|V(\mathcal{D})|\cdot Size_{\mathfrak{C}_{i},\mathcal{D}})$ for $\mathcal{P}_{i}$ being the one with the slowest running time of $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$, then
$\mathcal{O}\left(\sum_{i=1}^{r}{\tau(\mathfrak{C}_{i},\mathcal{D})}+|V(\mathcal{D})|\cdot\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}}\right)\leq\mathcal{O}(t_{D}^{r}).$
Theorem 4.6.
Before proving it, we will also state the following theorem, which is a formal reformulation of Theorem 1.3 from the introduction. The proof of Theorem 4.6 can be easily modified to work for Theorem 4.7.
Theorem 4.7.
Keep in mind that the following proof can be modified to work also for $DivMin$ as stated in Theorem 4.7. Instead of one diversity value $\ell$ for a tuple of partial solutions, we need to store a vector $L\in\{0,\ldots,d\}^{\binom{r}{2}}$ of pairwise Hamming distances between the solutions. Since for $S_{1},S_{2}\in S,\ P_{1},P_{2}\in P$, where $S$ and $P$ are disjoint, $HamDist(S_{1}\cup P_{1},S_{2}\cup P_{2})=HamDist(S_{1},S_{2})+HamDist(P_{1},P_{2})$, we can sum these vectors point-wise as we sum single diversity values. As the Hamming distance only grows, and we are interested in having the minimum distance of at least $d$, we can treat all the values greater than $d$ as equal to $d$. However, we cannot take advantage of keeping only the best diversity value for a tuple of entries because such vectors may be incomparable, so for one tuple of entries that represents multiple tuples of partial solutions, there may be up to $(d+1)^{\binom{r}{2}}$ different diversity vectors, which contributes to $d^{r^{2}}$ factor in Theorem 4.7.
Proof of Theorem 4.6.
  1. 1.
    Constructing the data corresponding to each core $\mathfrak{C}_{i}$.
    This step takes $\sum_{i=1}^{r}\tau(\mathfrak{C}_{i},\mathcal{D})$ time.
  2. 2.
    Constructing by induction for each $t\in V(\mathcal{D})$ set $\Pi_{Div}(t)$.
    • If $t=Leaf_{\mathcal{D}}(v)$,
      $\Pi_{Div}(t)=\{(w^{1},w^{2},\ldots,w^{r},\ell)|$ $\forall_{i\in[r]}w^{i}\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),$
      $\ell=f(\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))+(|V(G)|-1)f(0^{r})\}.$
    • Node $t$ has children $t_{1},\ldots,t_{\delta(t)}$.
      • If $\delta(t)=1$,
        $\Pi_{Div}(t)=\{(w^{1},w^{2},\ldots,w^{r},\ell)|\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1})}\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),\\ \ell=max\{\ell_{1}|\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1})}\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)\}\}.$
      • If $\delta(t)=2$,
        $\Pi_{Div}(t)=\{(w^{1},w^{2},\ldots,w^{r},\ell)|\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),(w^{1}_{2},\ldots w^{r}_{2},\ell_{2})\in\Pi_{Div}(t_{2})}\\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1},w^{i}_{2})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),\ell=max\{\ell_{1}+\ell_{2}-|V(G)|\cdot f(0^{r})|\\ \exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),(w^{1}_{2},\ldots w^{r}_{2},\ell_{2})\in\Pi_{Div}(t_{2})}s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1},w^{i}_{2})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)\}\}.$
      $((w^{1},w^{1}_{1},\ldots,w^{1}_{\delta(t)}),\ldots,(w^{r},w^{r}_{1},\ldots,w^{r}_{\delta(t)}))\in Process_{\mathfrak{C}_{1},\mathcal{D}}\times\ldots\times Process_{\mathfrak{C}_{r},\mathcal{D}}$ and checking whether for each $j\in[\delta(t)]$ there exists $(w^{1}_{j},\ldots w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$ for some $l_{j}\in\mathbb{N}$.
    Observe that by the definition of $\Pi_{Div}$, $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$ if and only if $w^{i}\in\Pi_{i}(t)$ for each $i\in[r]$, where $\Pi_{i}(t)$ is the set of $w\in\{0,1\}^{*}$ such that a $(\mathfrak{C}_{i},\mathcal{D}_{t},w)$-witness exists (see the proof of Theorem 4.4).
  3. 3.
    Checking whether there exists $(w^{1},w^{2},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$, such that $\ell\geq d$ and $w^{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ for each $i\in[r]$, where $q$ is the root of $\mathcal{D}$.
    This step takes $\mathcal{O}(\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}})$ time.
It remains to prove the correctness. Firstly, by Lemma 4.8, we will show that for each $S_{1},\ldots,S_{r}$ such that $(G,S_{i})\in\mathcal{P}_{i}$ there exists $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ for some $\ell\geq Div(S_{1},\ldots,S_{r})$. Secondly, by Lemma 4.9, we will show that for each $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ there exists $S_{1},\ldots,S_{r}$ such that $(G,S_{i})\in\mathcal{P}_{i}$ and ${Div(S_{1},\ldots,S_{r})=\ell}$.
@@ 4.1 Definitions
Proof.
@@
$=f((\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))+(|V(G)|-1)f(0^{r})=\ell.$
Node $\mathbf{t}$ has children $\mathbf{t_{1},\ldots t_{\delta(t)}}$. Let $j\in[\delta(t)]$ and $w^{1}_{j},\ldots,w^{r}_{j}\in\{0,1\}^{*}$. Assume that if for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t_{j}},w^{i}_{j})$-witness $\alpha_{i}^{j}$, then $(w^{1}_{j},\ldots,w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$ for some $l_{j}\geq Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{j}),\ldots,\allowbreak S_{\rho_{r}}(\mathcal{D}_{t_{j}},\alpha_{r}^{j}))$. Assume that for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness $\alpha_{i}$. For each $i\in[r]$, let $w^{i}_{j}\coloneqq\alpha_{i}(t_{j})$ and $\alpha_{i}^{j}(t^{\prime})\coloneqq\alpha_{i}(t^{\prime})$ for each $t^{\prime}\in V(\mathcal{D}_{t_{j}})$. Then observe that $\alpha_{i}^{j}$ is a properly defined $(\mathfrak{C}_{i},\mathcal{D}_{t_{j}},w^{i}_{j})$-witness. Hence, by the induction assumption, $(w^{1}_{j},\ldots,w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$ for some $l_{j}\geq Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{j}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{j}},\alpha_{r}^{j}))$. By the definition of a witness, $(w^{i},w^{i}_{1},\ldots,w^{i}_{\delta(t)})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)$ for each $i\in[r]$.
If $\delta(t)=1$, by the definition of $\Pi_{Div}$, ${(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)}$ for some $\ell\geq\ell_{1}$. Observe that for each $i\in[r]$ $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})$. So we get
@@
If $\delta(t)=2$, by the definition of $\Pi_{Div}$, ${(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)}$ for some $\ell\geq\ell_{1}+\ell_{2}-|V(G)|f(0^{r})$. Observe that for each $i\in[r]$, $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})\cup S_{\rho_{i}}(\mathcal{D}_{t_{2}},\alpha_{i}^{2}),$ where $S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1})\subseteq V(G_{t_{1}})$, $S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2})\subseteq V(G_{t_{2}})$ and $V(G_{t_{1}})\cap V(G_{t_{2}})=\emptyset$. So by Lemma 3.2
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))\\ =Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1}))+Div(S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2}))-|V(G)|f(0^{r})\\ \leq\ell_{1}+\ell_{2}-|V(G)|f(0^{r})\leq\ell.$
It follows that $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$ for some $\ell\geq Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))$. ∎ Observe that if $(G,S_{1},\ldots,S_{r})\in Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$, then for each $i\in[r]$ there exists a $(\mathfrak{C_{i}},\mathcal{D},w^{i})$-witness $\alpha_{i}$ for some $w^{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ such that $S_{i}=S_{\rho_{i}}(\mathcal{D},\alpha_{i})$. Then, by Lemma 4.8, $(w^{1},\ldots w^{r},\ell)\in\Pi_{Div}(t)$ for some $\ell\geq Div(S_{1},\ldots,S_{r})$.
Lemma 4.9.
Proof.
We use induction bottom-up on the decomposition tree.

### v2

@@ 3.1 Definitions
We start by computing the tables $Process_{\mathfrak{C}_{i},\mathcal{D}}$ and $Accept_{\mathfrak{C}_{i},\mathcal{D}}$ corresponding to each core in total time $\sum_{i=1}^{r}\tau(\mathfrak{C}_{i},\mathcal{D})$. Then, for each node $t$ of the decomposition, we construct by induction the set $\Pi_{Div}(t)$ of tuples $(w^{1},\ldots,w^{r},\ell)$ such that, for each $i$, there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness $\alpha_{i}$, and $\ell$ is the diversity of the list of partial solutions defined by this tuple of witnesses. If for some $(w^{1},\ldots,w^{r})$ there are multiple such tuples of witnesses, we keep only those that give the maximum value of the diversity. If $t$ is a leaf of $\mathcal{D}$, each witness $\alpha_{i}$ consists only of $w^{i}\in Process_{\mathfrak{C}_{i},\mathcal{D}}$. Observe that for any tuple $w=(w^{1},\ldots,w^{r})$, the only vertex whose membership in the list of partial solutions defined by this tuple can be nonzero is the vertex $v$ introduced at $t$. Moreover, we can compute the corresponding membership vector $m_{w}$ by simply applying the membership function. Then $\Pi_{Div}(t)$ consists of all such tuples $w$, together with the diversity value $\ell=f(m_{w})+(|V(G)|-1)f(0^{r})$. If $t$ has one child $t_{1}$, let $\Pi_{Div}(t)$ be the set of tuples $(w^{1},\ldots,w^{r},\ell)$ such that $\ell$ is the maximal value of $\ell_{1}$ among all $(w_{1}^{1},\ldots,w_{1}^{r},\ell_{1})\in\Pi_{Div}(t_{1})$ where each $(w^{i},w^{i}_{1})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)$. If $t$ has children $t_{1},t_{2}$, let $\Pi_{Div}(t)$ be the set of tuples $(w^{1},\ldots,w^{r},\ell)$ such that $\ell$ is the maximal value of $\ell_{1}+\ell_{2}-|V(G)|f(0^{r})$ among all pairs of $(w_{1}^{1},\ldots,w_{1}^{r},\ell_{1})\in\Pi_{Div}(t_{1}),(w_{2}^{1},\ldots,w_{2}^{r},\ell_{2})\in\Pi_{Div}(t_{2})$ where each $(w^{i},w^{i}_{1},w^{i}_{2})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)$. Note that we are using the monotonicity assumption in this argument. As we are joining pairs of partial solutions belonging to disjoint sets $V(G_{t_{1}}),V(G_{t_{2}})$, we can compute the diversity using Lemma˜3.8. Observe that for each node $t$ the set $\Pi_{Div}(t)$ can be computed in time $\prod_{i=1}^{r}|Process_{\mathfrak{C}_{i},\mathcal{D}}(t)|$. The last step is checking whether there exists $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ such that $\ell\geq d$ and each $w^{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ for $q$ being the root of $\mathcal{D}$. This can be done in $\prod_{i=1}^{r}|Accept_{\mathfrak{C}_{i},\mathcal{D}}|$ time.
We prove the correctness by combining the two following lemmas. Firstly, we show that for each $S_{1},\ldots,S_{r}$ such that $(G,S_{i})\in\mathcal{P}_{i}$ there exists $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ for some $\ell\geq Div(S_{1},\ldots,S_{r})$.
@@
Secondly, we show that for each $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ there exists $S_{1},\ldots,S_{r}$ such that $(G,S_{i})\in\mathcal{P}_{i}$ and $Div(S_{1},\ldots,S_{r})=\ell$}.
Lemma 3.10.
We prove Theorem˜3.6 before proving Lemma˜3.9 and Lemma˜3.10.
Proof of Theorem˜3.6.
  1. 1.
    First, constructing the data corresponding to each core $\mathfrak{C}_{i}$. This step takes $\sum_{i=1}^{r}\tau(\mathfrak{C}_{i},\mathcal{D})$ time.
  2. 2.
    Second, constructing by induction for each $t\in V(\mathcal{D})$ set $\Pi_{Div}(t)$.
    • If $t=Leaf_{\mathcal{D}}(v)$,
      $\Pi_{Div}(t)=\{(w^{1},\ldots,w^{r},\ell)$ $:\forall_{i\in[r]}w^{i}\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),$
      $\ell=f(\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))+(|V(G)|-1)f(0^{r})\}.$
    • If a node $t$ has children $t_{1},\ldots,t_{\delta(t)}$.
      • If $\delta(t)=1$,
        $\Pi_{Div}(t)=\{(w^{1},\ldots,w^{r},\ell):\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1})}\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),$
        $\ell=max\{\ell_{1}:\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1})}\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)\}\}.$
      • If $\delta(t)=2$,
        $\Pi_{Div}(t)=\{$ $(w^{1},\ldots,w^{r},\ell):\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),(w^{1}_{2},\ldots w^{r}_{2},\ell_{2})\in\Pi_{Div}(t_{2})}$
        $s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1},w^{i}_{2})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),\ell=max\{\ell_{1}+\ell_{2}-|V(G)|\cdot f(0^{r}):$
        $\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),(w^{1}_{2},\ldots w^{r}_{2},\ell_{2})\in\Pi_{Div}(t_{2})}\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1},w^{i}_{2})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)\}\}.$
      $((w^{1},w^{1}_{1},\ldots,w^{1}_{\delta(t)}),\ldots,(w^{r},w^{r}_{1},\ldots,w^{r}_{\delta(t)}))\in Process_{\mathfrak{C}_{1},\mathcal{D}}\times\cdots\times Process_{\mathfrak{C}_{r},\mathcal{D}}$
      and checking whether for each $j\in[\delta(t)]$ there exists $(w^{1}_{j},\ldots w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$ for some $\ell_{j}\in\mathbb{N}$.
    Observe that $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$ if and only if $w^{i}\in\Pi_{i}(t)$ for each $i\in[r]$, where $\Pi_{i}(t)$ is the set of $w\in\{0,1\}^{*}$ such that a $(\mathfrak{C}_{i},\mathcal{D}_{t},w)$-witness exists (see the proof of Theorem˜3.4).
  3. 3.
    The last step is checking whether there exists $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$, such that $\ell\geq d$ and $w^{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ for each $i\in[r]$, where $q$ is the root of $\mathcal{D}$, which takes $\mathcal{O}(\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}})$ time.
It remains to prove the correctness. Firstly, by Lemma˜3.9, we have that for each $S_{1},\ldots,S_{r}$ such that $(G,S_{i})\in\mathcal{P}_{i}$ there exists $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ for some $\ell\geq Div(S_{1},\ldots,S_{r})$. Secondly, by Lemma˜3.10, we have that for each $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ there exists $S_{1},\ldots,S_{r}$ such that $(G,S_{i})\in\mathcal{P}_{i}$ and $Div(S_{1},\ldots,S_{r})=\ell$}. As a result, we get the following corollary.
Corollary 3.11.
This finishes the proof of Theorem˜3.6. ∎
Proof of Lemma˜3.9.
@@
$=f(\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))+(|V(G)|-1)f(0^{r})=\ell.$
Node $\mathbf{t}$ has children $\mathbf{t_{1},\ldots t_{\delta(t)}}$. Assume that for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness $\alpha_{i}$ and let $w^{i}=\alpha_{i}(t)$. Let $t_{j}$ be a child of $t$. For $i\in[r]$ let $w_{j}^{i}\coloneq\alpha_{i}(t_{j})$. Observe that $\alpha_{i}^{j}\coloneq\alpha_{i}|_{V(\mathcal{D}_{t_{j}})}$ is a $(\mathfrak{C}_{i},\mathcal{D}_{t_{j}},w^{i}_{j})$-witness. Then, by the induction assumption, $(w^{1}_{j},\ldots,w^{r}_{j},l_{j})\in\Pi_{Div}(t_{j})$ for some $l_{j}\geq Div(S_{\rho_{1}}(\mathcal{D}_{t_{j}},\alpha_{1}^{j}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{j}},\alpha_{r}^{j}))$. Also note that for each $i\in[r]$, $(w^{i},w^{i}_{1},\ldots,w^{i}_{\delta(t)})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)$, as $\alpha_{i}$ is a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness.
If $\delta(t)=1$, by the definition of $\Pi_{Div}$, $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$} for some $\ell\geq\ell_{1}$. As for each $i\in[r]$ $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})$, we get
@@
If $\delta(t)=2$, by the definition of $\Pi_{Div}$, $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$} for some $\ell\geq\ell_{1}+\ell_{2}-|V(G)|f(0^{r})$. Observe that for each $i\in[r]$, $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})\cup S_{\rho_{i}}(\mathcal{D}_{t_{2}},\alpha_{i}^{2}),$ where $S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1})\subseteq V(G_{t_{1}})$, $S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2})\subseteq V(G_{t_{2}})$ and $V(G_{t_{1}})\cap V(G_{t_{2}})=\emptyset$. So by Lemma˜3.8
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))$
$=Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1}))+Div(S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2}))-|V(G)|f(0^{r})$
$\leq\ell_{1}+\ell_{2}-|V(G)|f(0^{r})\leq\ell.$
It follows that $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$ for some $\ell\geq Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))$}. ∎
Proof of Lemma˜3.10.
Again we use induction bottom-up on the decomposition tree.